# Jeffrey Bub

### v2

@@ 2 The Measurement Problem
There are different formulations of the measurement problem, but the classic formulation is von Neumann’s statement in [46, p. 351]. Von Neumann points to two fundamentally different ways in which the state of a quantum system can change: the change that occurs when a system undergoes measurement (he calls this Process 1), and the change when a system evolves dynamically under the action of an energy operator $H$ without measurement (Process 2). When a system evolves dynamically without being measured, the state $|\psi\rangle$ undergoes a continuous, reversible, unitary transformation in time $t$ to the state
@@
But when a system is measured, the state $|\psi\rangle$ undergoes a discontinuous, irreversible, non-unitary transformation to one of the eigenstates $|\phi_{n}\rangle$ of the measured observable with probability $|\langle\phi_{n}|\psi\rangle|^{2}$. That is, the transition is from the pure state $W=P_{|\psi\rangle}$ to a mixed state
@@
Consider, for example, a macrosystem composed of $N$ elementary 2-state systems. The Hilbert space has $2^{N}$ dimensions, so as $N\rightarrow\infty$ the dimensionality becomes uncountable, $2^{N}\rightarrow 2^{\aleph_{0}}=\aleph_{1}$. A Hilbert space with an uncountable number of dimensions is non-separable and breaks down into orthogonal sectors, each of which is a separable Hilbert space spanned by a countable basis of vector states. The sectors are equivalence classes of vectors, where $|\Psi\rangle=\otimes_{i}|\psi_{i}\rangle$ and $|\Phi\rangle=\otimes_{i}|\phi_{i}\rangle$ are equivalent if and only if $\sum_{i}|\langle\phi_{i}|\psi_{i}\rangle-1|$ converges. Macro-observables and macrostates can be defined such that a macro-observable takes the same eigenvalue for all the vectors in a sector. Dynamical evolutions in which only a finite number of elementary systems undergo a change of state leave the macrostate in the same sector. The whole Hilbert space is a direct sum of these orthogonal sectors, so the transition to infinity automatically yields a superselection rule, a restriction on the superposition principle.
@@
The second phase involves a macroscopic detector, $D$, such as an avalanche photodiode (a type of photodetector), in which an exponentially increasing number of elementary systems becomes entangled with the $S$ and $M$ systems in a chain reaction that is not reversible. The density matrix of the composite system becomes almost block-diagonal for large $N$ and, as $N\rightarrow\infty$, the off-diagonal terms vanish in the limit and there is zero interference between sectors. At the infinite limit the vector state of the composite system, a sum of terms from different sectors, represents a mixture of composite system states—the density operator is a sum of projection operators onto product states of $S,M,D$, with coefficients representing the probabilities of these states:
@@
To convey an idea of how sectorization works, Van Den Bossche and Grangier consider a simple example of NN spin-$1/2$ systems. The spins could be positive or negative in the $z$-direction, $|\!\uparrow_{i}\rangle,|\!\downarrow_{i}\rangle$, or positive or negative in some other direction, say the $x$-direction, $|+_{i}\rangle,|-_{i}\rangle$. If all the spins are aligned in the positive $z$-direction, the state of the composite system is
$|\Psi_{N}\rangle=\otimes_{i=1}^{N}|\!\uparrow_{i}\rangle$ (7)
In the limit $N\rightarrow\infty$ this state ends up in a particular sector of Hilbert space. States of the composite system with all but a finite number of elementary spin states in the positive $z$-direction end up in the same sector and are non-orthogonal (unless the scalar product of two component elementary states is zero for some $i$). If all the spins are aligned in the positive $x$-direction, the state of the composite system is
$|\Phi_{N}\rangle=\otimes_{i=1}^{N}|+_{i}\rangle$ (8)
The scalar product of $|\Psi_{N}\rangle$ and $|\Phi_{N}\rangle$ is
$\langle\Psi_{N}|\Phi_{N}\rangle=(1/\sqrt{2})^{N}$ (9)
which tends to 0 as $N\rightarrow\infty$. So in the limit these two states end up in different sectors.
What’s new about the limiting case is that some vector states represent mixtures. At the infinite limit, because there is no interference between states belonging to different sectors, states represented by superpositions of vectors from different sectors represent mixtures. Mixtures of states from the same sector are not represented by vector states.
These considerations show that the measurement problem is resolved at the infinite limit. But infinity never comes. For macrosystems with very large but finite $N$, we can say only that interference between different macrostates is very small, less than any small number $\epsilon(N)$, and gets smaller and smaller, and harder and harder to detect, as the composite system includes more and more elementary systems. The approximate sectorization of Hilbert space as the number of elementary particles in a system increases without limit, and the consequent negligible interference between different macrostates, explains the classicality of the macroworld for all practical purposes—FAPP to use Bell’s acronym—and we can see why we experience the macroworld as classical or Boolean without the usual appeal to decoherence via loss of information to the environment. This is a nice result, but it does not explain the occurrence of a definite outcome in a quantum measurement process, so it does not explain how von Neumann’s Process 1 is reducible to Process 2.
@@ Acknowledgements
Thanks to Michel Janssen, Allen Stairs, and especially Michael Cuffaro and Jeremy Butterfield, for insightful and extremely helpful comments.

### v3

@@ 2 The Measurement Problem
There are different formulations of the measurement problem, but the classic formulation is von Neumann’s statement in [47, p. 351]. Von Neumann points to two fundamentally different ways in which the state of a quantum system can change: the change that occurs when a system undergoes measurement (he calls this Process 1), and the change when a system evolves dynamically under the action of an energy operator $H$ without measurement (Process 2). When a system evolves dynamically without being measured, the state $|\psi\rangle$ undergoes a continuous, reversible, unitary transformation in time $t$ to the state
@@
But when a system is measured, the state $|\psi\rangle$ undergoes a discontinuous, irreversible, non-unitary transformation to one of the eigenstates $|\phi_{n}\rangle$ of the measured observable with probability $|\langle\phi_{n}|\psi\rangle|^{2}$. That is, the transition (for a non-selective measurement) is from the pure state $W=P_{|\psi\rangle}$ to a mixed state
@@
Consider, for example, a macrosystem composed of $N$ elementary 2-state systems, say spin-1/2 systems. The Hilbert space representation of the spin algebra for finite $N$ is irreducible: it contains no proper subspace whose vectors remain in the subspace after being acted on by a spin operator, and all representations of the spin algebra are unitarily equivalent to the Pauli representation (von Neumann’s uniqueness theorem [43]). As $N\rightarrow\infty$ the number of dimensions $2^{N}$ becomes uncountable: $2^{N}\rightarrow 2^{\aleph_{0}}=\mathfrak{c}=|\mathbb{R}|$ and the uniqueness theorem no longer holds. [^4] Thanks to Krzysztof Sienicki for pointing out a correction to my formulation of this point in an an earlier version. A Hilbert space with an uncountable number of dimensions is non-separable and decomposes into into a direct sum of $\infty$-dimensional Hilbert spaces, each of which provides a different irreducible representation of the algebra of observables. These inequivalent representations correspond to different macrostates of the system, where a macrostate is identified with an equivalence class or sector of microstates. A macro-observable takes the same value for every microstate in a given sector. For example, the spin density of an infinite array of spin-1/2 systems defines a macro-observable, the polarization of the macrosystem, that takes different values for different sectors. Dynamical evolutions in which only a finite number of elementary systems undergo a change of state leave the macrostate in the same sector. This sectorization of Hilbert space associated with inequivalent representations of the spin algebra yields a restriction on the superposition principle, a decomposition of the Hilbert space into superselection sectors.
@@
The second phase involves a macroscopic detector, $D$, such as an avalanche photodiode (a type of photodetector), in which an exponentially increasing number of elementary systems becomes entangled with the $S$ and $M$ systems in a chain reaction that is not reversible. The density matrix of the composite system becomes almost block-diagonal for large $N$ and, as $N\rightarrow\infty$, the off-diagonal terms vanish in the limit and there is zero interference between sectors. At the infinite limit the vector state of the composite system, a sum of terms from different sectors, represents a mixture—the density operator can be represented as a sum of projection operators onto composite system states, with coefficients representing the probabilities of these states:
@@
What’s new about the transition to infinity is that vector states need not be pure states on the algebra generated by sectorization. So, in effect, at the infinite limit, because there is no interference between states belonging to different sectors, states represented by superpositions of vectors from different sectors represent mixtures. Vector states representing superpositions of states from the same sector do not represent mixtures.
These considerations show that the measurement problem is resolved at the infinite limit. But infinity never comes. For macrosystems with very large but finite $N$, we can say only that interference between different macrostates is very small, less than any small number $\epsilon(N)$, and gets smaller and smaller, and harder and harder to detect, as the composite system includes more and more elementary systems. The approximate sectorization of Hilbert space as the number of elementary particles in a system increases without limit, and the consequent negligible interference between different macrostates, explains the classicality of the macroworld for all practical purposes—FAPP to use Bell’s acronym—and we can see why we experience the macroworld as classical or Boolean. This is informative, but it does not explain the occurrence of a definite outcome in a quantum measurement process, so it does not explain how von Neumann’s Process 1 is reducible to Process 2.
@@ Acknowledgements
Thanks to Michel Janssen, Allen Stairs, and especially Michael Cuffaro and Jeremy Butterfield, for insightful and extremely helpful comments. Many thanks also to Krzysztof Sienicki who suggested several edits in a technical comment in arXiv: 2512.24198 on an earlier version of this paper.